OpenStudy (laddiusmaximus):

using galileo's equations of motion which are given by s(t)=s0+v0t-(1/2)gt^(2) and v(t) = (ds/dt)=v0-gt g=9.8m/s^(2) for gravity a ball tossed in the air vertically from ground level returns to earth 5 seconds later. Find the initial velocity and the maximum height of the ball

OpenStudy (laddiusmaximus):

maximum height would be at 2.5 seconds right?

OpenStudy (laddiusmaximus):

so for intial veloctiy that is 9.3 (2.5) =24.5 correct?

OpenStudy (anonymous):

Calculate the maximum; the vertex to find max height

OpenStudy (laddiusmaximus):

9.8(2.5)

OpenStudy (anonymous):

yes max would be half way at 2.5 seconds

OpenStudy (anonymous):

and since it is thrown from the ground, \[s_0=0\] because that is the inital height

OpenStudy (anonymous):

and you know \[v_0t-4.9t^2=0\] when \[t=5\] so solve for \[v_0\] in \[v_0\times 5-4.9(5)^2=0\]